Janice can mow the lawn in 2 hours. her friend can mow the same lawn in 2 hours. How long would it take them to mow the lawn if they had two lawn mowers and worked for the same length of time?
Answer by checkley77(12844)

1/2+1/2=1/X ADD THE FRACTIONS ON THE LEFT SIDE.
1=1/X CROSS MULTIPLY.
X=1 HOUR WORKING TOGETHER.